### v1

# microtypeUnable to apply patch

Abstract
Empathetic response generation is a desirable aspect of conversational agents, crucial for facilitating engaging and emotionally intelligent multi-turn conversations between humans and machines. Leveraging large language models for this task has shown promising results, yet challenges persist in ensuring both the empathetic quality of the responses and retention of the generalization performance of the models. In this paper, we propose a novel approach where we construct theory-driven preference datasets and use them to align LLMs with preference optimization algorithms to address these challenges. To measure empathetic response generation, we employ the EmpatheticDialogues dataset, assessing empathy with the diff-Epitome and BERTscore metrics, and evaluate the generalization performance on the MMLU benchmark. We make all datasets, source code, and models publicly available.111github.com/ondrejsotolar/empo
1 Introduction
Empathetic response generation (ERG) focuses on tuning a conversational agent toward understanding the user’s situation, feelings, and experience to generate appropriate, human-like responses. This goal was first attempted by the earliest rule-based chatbots like ELIZA (Weizenbaum, 1966), then more recently with deep learning approaches such as BlenderBot (Roller et al., 2021; Xu et al., 2022; Komeili et al., 2022; Shuster et al., 2022). The introduction of instruction-tuned large language models (LLMs) such as ChatGPT OpenAI (2023) has challenged the contemporary approaches to ERG. These models can participate in multi-turn conversations without additional training and are implicitly capable of generating empathetic conversations (Lee et al., 2022a).
LLMs also have shown tremendous generalization capabilities and can draw on world knowledge obtained by training on Internet-size data. These traits were not present in the earlier non-LLM conversational agents. Recent developments focus on approaches that keep the generalization abilities of LLMs while improving empathy in responses, including prompt-engineering methods that do not change the model’s weights but rather tune the prompt content to assist the generation by in-context learning (Wang et al., 2024), often assisted by retrieval (Qian et al., 2023).
In the current study, we propose a data-driven solution for ERG with LLMs, based on preference optimization: aligning LLMs via preference optimization algorithms. First, we build a preference dataset using the benchmark EmpatheticDialogues dataset (Rashkin et al., 2019). The dataset contains short multi-turn human-to-human dialogues grounded by emotion labels. We leverage this property to extract responses from the corpus of the polar opposite emotion label using Plutchik’s wheel (Plutchik, 2001) such that each prompt is paired with preferred and non-preferred completions. We then fine-tune a foundational LLM using Direct Preference Optimization (Rafailov et al., 2024) to generate responses aligned with the preferred candidate response.
Using diff-Epitome (Lee et al., 2022b), a model-based empathy metric for multi-turn conversations, we show that training LLMs with our preference dataset improves ERG. We also investigate the models’ general language understanding using the MMLU benchmark to show how our method impacts the performance on other tasks. Our method is analogous to providing guardrails with helpful/harmful preference datasets (Bai et al., 2022). Contemporary prompt-engineering methods or additional training can be applied to models aligned in this way to adapt them further for any task. We also share novel observations from searching over the hyperparameter configuration space and provide code to apply our method to other datasets and models.
2 Related Work
Systems specialized for ERG such as KEMP (Li et al., 2022), CEM (Sahand Sabour, 2021), MIME (Majumder et al., 2020), EmpDG (Li et al., 2020), and DIFFUSEMP (Bi et al., 2023) or universal conversational agent like Blenderbot were not built for general instruction-following. Recent approaches such as (Wang et al., 2024; Qian et al., 2023; Li et al., 2024) employ prompt engineering with LLMs to some success. However, their approach differs principally from ours; we restructure the problem of ERG as an LLM alignment problem and produce a method for creating preference datasets and a modeling approach rather than a prompt construction algorithm.
To our knowledge, the current study is the first to explore aligning LLMs for ERG via preference optimization.
Datasets
Several dialogue datasets, including IEMOCAP (Busso et al., 2008), MELD (Poria et al., 2019), DailyDialog (Li et al., 2017), Herzig et al., 2016, EmotionLines (Hsu et al., 2018), EmpathicReactions (Buechel et al., 2018), and EmoContext (Chatterjee et al., 2019), contain emotion labels. However, these datasets are either labeled with only a small set of emotions, limited by size, or lack a multi-turn character.
The EmpatheticDialogues (ED) is a dataset with 25K human-to-human dialogues and 32 emotion labels, which are derived from biological responses (Ekman, 1992; Plutchik, 1980) to larger sets of subtle emotions derived from contextual situations (Skerry and Saxe, 2015). It is the benchmark dataset on empathetic conversation as it addresses preceding datasets’ limitations.
Earlier deep learning-based conversational agents required large datasets for training from scratch, often at the expense of quality. A subset of ED was analyzed for listener-specific response intents, creating the EDOS dataset (Welivita et al., 2021), which includes 1M movie dialogues annotated with a BERT-based classifier. Recently, synthetic datasets like Chinese SMILE (Qiu et al., 2024) and Korean SoulChat (Chen et al., 2023) have been generated using LLMs.
However, ED and its successors have been criticized for their data annotation and model evaluation approaches (Debnath and Conlan, 2023). In treating ERG as an LLM alignment task, our approach addresses some of these drawbacks.
Evaluating ERG
Evaluation of empathy in dialogues has primarily focused on human evaluation and lexical overlap/semantic similarity with reference dialogues. The former is limited by the highly subjective nature of the task, the need to train annotators, and, usually, small sample sizes. We have found the latter to have little relevance for measuring empathy, given the highly creative nature of generative LLM responses, in agreement with the findings of Liu et al., 2016. We instead rely on model-based empathy metrics, which use classification models trained specifically for the task of measuring empathy in multi-turn conversations: diff-EPITOME (Lee et al., 2022b) which is based on the EPITOME (Sharma et al., 2020) classifiers. For measuring the generalization abilities, we use the standard benchmarks such as MMLU (Hendrycks et al., 2020), using the lm-evaluation-harness (Bommasani et al., 2023).
LLM Alignment by Preference Optimization
Aligning generative models with human feedback has improved their helpfulness, factual accuracy, and ethical behavior, among other aspects Ouyang et al. (2022). Methods like RLHF (Christiano et al., 2017), including training algorithms such as PPO (Schulman et al., 2017) and DPO (Rafailov et al., 2023) have consistently been more effective than relying solely on supervised fine-tuning (SFT). Human feedback can take many forms: PPO requires human preferences to rank the generations, and DPO requires datasets of prompts and pairs of preferred/rejected completions.
3 Experimental Setup
We conduct our experiments using the Zephyr-7B (Tunstall et al., 2023) model, a variant of the Mistral-7B foundational LLM (Jiang et al., 2023) fine-tuned for multi-turn dialogues using the UltraChat dataset (Ding et al., 2023). The smallest variant with 7 billion parameters already provides good general language understanding ability as it matches LLama-2-70B on many NLP benchmarks (Beeching et al., 2023). For all training steps, we explore the hyperparameter configuration space and optimize for a minimal impact on generalization while improving empathy scores.
Supervised Fine-tuning
As Tunstall et al., 2023 showed, SFT is a necessary first step in alignment, which ensures that the preference dataset is in-domain for the aligned model. We perform SFT using the standard causal auto-regressive objective using the individual dialogues from EmpatheticDialogues. As per the definition of the dataset, the odd turns are considered the "user prompts" and the even turns the "responses". We limit the SFT training to the even-indexed turns by masking the odd turns to ignore them in the loss-function computation. For computational efficiency, we fine-tune using LoRA adapters (Hu et al., 2021).
Preference Optimization
For DPO , we build a preference dataset from the EmpatheticDialogues consisting of preferred/rejected completions. For each dialogue, we target the last even turn (the last "response" to user) as the generation target while including the previous turns as context. This is the standard way of processing the dataset, also done in previous works.
For constructing the completion pairs, we leverage a property of the ED dataset: each dialogue is associated with an emotion label. These were used as grounding during the ideation of each dialogue. For the preferred completion, we take the ground truth - the original response. For its rejected counterpart, we use Plutchik’s wheel of emotions (Figure 4), and the derivative emotional dyads (Figure  5,  Plutchik, 2001) to find the polar opposite emotion labels resulting in a lookup table (see A.3). For each completion, we randomly select one from the group of completions labeled with this opposite label. Because the stability can suffer from this random selection, we draw a fresh random completion for each new epoch of training and search for the hyperparameter configuration to offset the repetition of the preferred completions.
Model MMLU (5s) $\uparrow$ diff-ER $\downarrow$ diff-EX $\downarrow$ diff-IP $\downarrow$ FBert $\uparrow$
[…]
Table 1: Language understanding, empathy, and semantic similarity with EmpatheticDialogues. All presented results are means ($\pm$ SD) of scores over multiple (4) training runs with the same hyperparameters but different seeds.
Empathy Evaluation
For measuring empathy, we use the diff-Epitome metric Lee et al. (2022a): a model-based metric specifically developed for measuring empathy in dialogues. diff-Epitome is an evolution of the Epitome classifiers (Sharma et al., 2020), which measure empathy in dialogue on a scale (0-2) from none (0), through weak (1), to strong (2) on three dimensions: empathetic responses (ER), explanations (EX), and interpretations (IP). The diff-Epitome uses a similar but open-ended continuous scale (0-) on the same dimensions by averaging the scores across the entire dataset, thus providing a measure of difference from the ground truth. Given our data-driven approach, we consider a lower difference from the ground truth better.
In preliminary human-evaluation experiments with lexical-overlap metrics such as BLEU or ROUGE and vector-similarity such as BERTscore, we found them uncorrelated with the perceived quality of the generated responses. This was especially evident with the largest LLMs, such as GPT4 and Claude, which generated high-quality responses yet received low scores from overlap metrics. We attribute this to the highly diverse nature of LLM generations. However, we keep using the semantic similarity BERTscore for sanity and increased metric diversity.
Language Understanding Evaluation
We measure the general language of the models with established benchmarks, such as the Massive Multitask Language Understanding (MMLU) Hendrycks et al. (2020) MMLU is designed to assess how LLMs learn and apply knowledge across a variety of domains by measuring their performance on 57 different multiple-choice style question tasks from STEM, humanities, and world knowledge. This benchmark requires the model to be capable of general instruction following.
4 Results
Table 1 compares the performance of the baseline, unaligned model to the supervised fine-tuned model (SFT) and to the SFT model further aligned with a preference optimization algorithm (DPO).
SFT
We observed a clear improvement across the board in empathy and similarity metrics provided by the SFT step. However, the more intensely the model is fine-tuned, the more it over-fits on the ED dataset, and the general performance drops, as shown with language understanding metrics in Figure 1. Mitigating this, but not avoiding it, is possible through careful hyperparameter optimization. First, as Tunstall et al., 2023 suggests, we limited the training to one epoch. The key to reducing over-fitting was using a large lora_rank ($r$) and low lora_alpha ($\alpha$) relative to the rank. The large $\alpha$ supports retention of the original model’s abilities, and the rank $r$ controls the impact size of the adapter’s fine-tuned weights. Figure 2 shows the empathy metrics saturating with $\alpha\approx\frac{r}{4}$. Furthermore, we observed that increasing the rank above 1024 brings diminishing returns: the computational efficiency of LoRA vanishes while the over-fitting problem remains. We use the best configuration jointly for empathy and generalization ($r\!=\!1024,\alpha\!=\!256,lr\!=\!1e\!-\!5$) to report SFT results and to perform DPO experiments.
Figure 1: SFT: Impact of LoRA $\alpha$ on the MMLU score. Trained with: learning_rate=1-e5, batch_size=64.
Figure 2: SFT: Impact of LoRA adapter rank on the diff-Epitome score during SFT. Trained with: learning_rate=1-e5, batch_size=64.
DPO
The preference alignment with DPO improved the empathy metrics over the SFT while retaining the model’s general performance. Nevertheless, its hyperparameters need to be set suitably to achieve this. Unsuitable hyperparameter configuration leads to training instability in addition to over-fitting (see Figure 3). Notably, we faced problems with the stability of the training introduced by the dataset construction process: random selection of rejected completions from the group of dialogues labeled with polar opposite emotions. We solved it by training for more epochs: re-drawing new randomly selected rejected completions for the same preferred ones for each epoch. The optimum configuration for high empathy scores while controlling over-fitting was training for three epochs, same as in Tunstall et al., 2023, with low $\beta$ ($0.025$) but a relatively high learning rate ($51e\!-\!$ 5). We did not optimize BERTscore in SFT or DPO training, and results indicate no correlation with the diff-Epitome metric.
Figure 3: Stability of the DPO training measured by the spread of metric scores between multiple runs with the same hyperparameters. A smaller spread signifies higher training stability. We measure the mean and spread of the diff-Epitome metric, averaged across its three dimensions (ER, EX, IP: lower is better). We observe that setting $\beta\!>>\!0.05$ leads to over-fitting on the preferred completions from the training set.
5 Conclusion
We introduced a new method for constructing preference datasets to align large language models (LLMs) for empathetic response generation while maintaining their overall language understanding capabilities. We validate our findings using the EmpatheticDialogues dataset and evaluate using the diff-Epitome and BERTscore metrics for empathy and standard language understanding benchmarks. Our method achieves the desired alignment of LLMs and provides a robust foundation for further enhancements, such as through prompt engineering. Furthermore, our method is directly applicable to other datasets with emotion labels.
Limitations
Empathetic Responses
Lahnala et al., 2022 argue that most NLP research defines empathy loosely as understanding and appropriately responding to others’ emotions, focusing mainly on detecting sentiment, emotions, or supportive interactions in text as indicators of empathy. This perspective assumes that systems achieving emotional recognition and responding in line with the target’s sentiment are empathetic. However, this approach overlooks the critical aspect of cognitive empathy, which involves understanding another person’s perspective, a gap highlighted by established human empathy theories Debnath and Conlan (2023).
Instead, we use a fully data-driven approach. We utilize the EmpatheticDialogs dataset as a source of the ground-truth empathetic responses. Even though there is criticism (Debnath and Conlan, 2023), dialogues in this dataset are written by human conversation partners and thus align well with real-world empathetic interactions.
Appendix A Training Details
A.1 Model Settings
We trained the models using the HuggingFace Transformers library. In the SFT step, we applied standard sequence-to-sequence training with cross-entropy loss on tokens from the user prompts (not on the tokens from "replies"). In both the SFT and DPO training steps, we optimized the model with AdamW optimizer and effective batch size of 64. We used learning rate of 1e-5 without warmup and a cosine lr decay. The models were trained in bf16 precision.
A.2 Hardware
To train our models, we used two NVIDIA A100 80GB GPUs. We trained LoRA adapters with ranks ranging from 16 to 2048 for models with 7BM parameters. The total training wall time, including preliminary experiments, was 17 days.
A.3 DPO Preference Dataset
For training with DPO, we introduced a method to construct a preference dataset from the EmpatheticDialogues dataset. With each training run or a single epoch within a multi-epoch run, the dataset is created anew as described in Section 3. This section contains auxiliary material to support the description of the preference dataset creation.
Each dialog in the EmpatheticDialogues dataset is associated with an emotion label. To construct the training example for DPO, we paired the preferred dialog completion (in our case, using the ground truth) with one rejected competition. To find the rejected completion, we proposed using opposite emotion labels. The opposites are based on two theory-based sources: Plutchik’s wheel (Figure 4), emotional dyads (Figure 5), and we proposed the rest ourselves. The opposites form the lookup table 2, which is queried each time a preferred/rejected pair is constructed.
Figure 4: Plutchik’s wheel of emotions (Creative Commons). It shows eight basic emotions: joy, trust, fear, surprise, sadness, anticipation, anger, and disgust. The wheel of emotions groups these eight basic emotions based on the physiological purpose of each into polar coordinates reflecting their similarity and intensity.
Figure 5: Plutchik’s emotion’s dyads (Creative Commons). When two emotions are elicited together, they form the primary dyad. If they are one petal apart, it is a secondary dyad; if they are two petals distant from each other, it is a tertiary dyad. Opposite dyads are on the opposite side.
Emotion label Opposite label Source
afraid angry wheel
angry afraid wheel
sad joyful wheel
grateful disgusted wheel
surprised anticipating wheel
trusting disgusted wheel
disgusted trusting wheel
anticipating surprised wheel
content anxious wheel
apprehensive annoyed wheel
joyful sad wheel
proud ashamed dyads
prepared anxious dyads
ashamed proud dyads
guilty proud dyads
nostalgic hopeful dyads
anxious content dyads
hopeful nostalgic dyads
sentimental apprehensive
jealous faithful
embarrassed confident
excited devastated
annoyed apprehensive
lonely caring
faithful jealous
terrified furious
confident embarrassed
furious terrified
disappointed impressed
caring lonely
impressed disappointed
devastated excited
Table 2: The opposite emotion labels lookup table. The emotion labels are sourced from the EmpatheticDialogues dataset. Each pair of opposites is associated with how the pair was determined: using Plutchiks’s wheel, emotion dyads, or our proposal.